OpenStudy (anonymous):

determine the slope and y intercept of the line given the equation : 2x-5y=45

OpenStudy (anonymous):

2x-45 = 5y 2/5x - 9 = y

OpenStudy (anonymous):

To find the slope, rearrange your equation so it matches y=mx+c. To find where the line cuts the y-axis let x equal to zero and sub it into your equation.

OpenStudy (anonymous):

for the slope i got 2/5 and intercept i got (0,-9)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

2x-45 = 5y 2/5x - 9 = y

OpenStudy (anonymous):

kay- thanks :)

OpenStudy (anonymous):

Perfectly correct.

OpenStudy (anonymous):

yaeh! :) lol